# Emergent Fracton Hydrodynamics in the Fractional Quantum Hall Regime of Ultracold Atoms

### v1

@@ I Introduction
Recent experimental and technical advances enabled the study of quantum Hall physics with synthetic quantum matter. Artificial gauge fields have been realized for ultracold quantum gases in the continuum by using the equivalence of neutral particles under rotation and charged particles in a magnetic field [1, 2, 3, 4, 5]. By properly designing the rotating trapping potential, recent work has demonstrated the geometric squeezing of a condensate into the Lowest Landau Level (LLL) [6, 7, 8, 9]. Artificial magnetic fields have also been realized for ultracold quantum gases by spatially dependent optical couplings [10, 11, 12] and for charge neutral atoms in optical lattices by laser dressing [13, 14, 15, 16, 17, 18, 19, 20, 21]. The latter approach enabled the observation of interaction-induced propagation of chiral excitations in the few-body limit of fractional quantum Hall (FQH) states [22, 23]. Nonetheless, adiabatically preparing fractional FQH ground states in the many-body limit and probing their intricate properties remain central open challenges.
@@
In this work, we propose to identify key signatures of FQH physics in the far-from-equilibrium relaxation dynamics instead. Recently, it has been shown that quenching the effective mass tensor excites the dipole and graviton modes of FQH states [24, 25, 26]. These quench protocols are initialized with the ground state of the system. By contrast, here, we will avoid the challenge of preparing the interacting ground state and explore the unique signatures in the the far-from equilibrium dynamics of a high temperature state in the LLL. Interacting particles in the LLL are described by effective one-dimensional models which conserve both the global charge as well as the associated center of mass (or equivalently the dipole moment) [27, 28, 29, 30, 31, 32, 33, 34, 35, 36, 30, 37]. Due to these effective conservation laws, the mobility of the particles is constrained. Following the predictions of fracton hydrodynamics, that emerges in systems with both charge and dipole conservation [38, 39, 40], we show that these systems exhibit slow relaxation dynamics due to interaction-induced intra-Landau-level scattering processes. This unconventional relaxation dynamics thus probes and characterizes the FQH physics of the LLL.
@@
Our results are organized as follows. In Sec. II we introduce the effective model for the LLL and in Sec. III we analyze the emergent fracton hydrodynamics of interacting particles in the LLL; we find anomalous relaxation dynamics of the density-density autocorrelation function as well as the bipartite fluctuations of the particle number in a subregion of the system. In Sec. IV we discuss experimental platforms for realizing FQH physics with ultracold atoms, focusing both on rotating quantum gases and on laser-dressed atoms in optical lattices. In both cases, we analyze the full, unprojected dynamics and determine the conditions for conserving the dipole moment in real space. An outlook and discussions are provided in Sec. V.
@@ II Effective model for the Lowest Landau Level
Figure 1: Projection onto the lowest Landau level. (a) Illustration of the continuum model on a cylinder. (b) Illustration of the Harper-Hofstadter model on a torus in the Landau gauge; flat bands are realized for specific lattice geometries, as discussed in the main text. (c) When projecting onto the lowest band, the contact interaction gives rise to center-of-mass (i.e., dipole-moment) conserving hopping terms, whose matrix elements decay with typical length scale $\sigma$. (d) For the continuum model the typical length scale $\sigma$ depends only on the ratio $L_{y}/l_{b}$. (e) For the Harper-Hofstadter model $\sigma$ can be tuned by changing the hopping ratio $t_{y}/t_{x}$ ($\sigma$ is numerically evaluated for system size of $N_{x}=44$ and $N_{y}=22$).
@@
In the continuum, the single-particle Hamiltonian that describes the system is $H_{0}=\frac{1}{2m}(\mathbb{p}-\mathbb{A})^{2}$, where $(\mathbb{p}-\mathbb{A})$ is the generalized momentum in the presence of the magnetic field, and $\mathbb{A}=B\,(0,x)$ is the vector potential in the Landau gauge. The properties of the system are determined by the magnetic length $l_{b}=\sqrt{\frac{\hbar}{eB}}$, the number of magnetic fluxes $N_{\phi}=\frac{L_{x}L_{y}}{2\pi l^{2}_{b}}$, and the cyclotron frequency $\omega_{c}=\frac{eB}{m}$. Assuming an infinite cylinder geometry with periodic boundary conditions in the $y$-direction, Fig. 1(a), the basis that diagonalizes the single-particle Hamiltonian is [27, 28, 29, 36], $\psi_{n}^{\gamma}(\mathbb{r})=\frac{1}{\sqrt{{\pi}n!2^{n}L_{y}}}H_{n}\bigg{(}\frac{x-x_{\gamma}}{l_{B}}\bigg{)}e^{-\frac{1}{2}(\frac{x-x_{\gamma}}{l_{b}})^{2}}e^{ik_{\gamma}y}$, where $H_{n}$ are the Hermite polynomials, $x_{\gamma}=2\pi\gamma\frac{l^{2}_{b}}{L_{y}}$ with $\gamma\in[0,N_{\phi}-1]$ determine the center of the wave function, and $k_{\gamma}=\frac{2\pi\gamma}{L_{y}}$ is the momentum in $y$-direction. The associated single-particle eigenvalues are $E_{n}^{\gamma}=\hbar\omega_{c}(n+\frac{1}{2})$. Assuming that particles initially occupy only one Landau level, density-density interactions $V=\sum_{i,j}V_{i,j}n_{i}n_{j}$ of magnitude much smaller than the inter-level spacing $\omega_{c}$, give rise to an effective dynamics that is confined to the same Landau level. The Hamiltonian is obtained by projecting the interacting Hamiltonian $H=H_{0}+V$ onto the Landau level [36]
$P_{0}HP_{0}=\sum_{q=0}^{N_{\phi}-1}\sum_{p=0}^{\frac{N_{\phi}}{2}-1}\sum_{k=p+1}^{\frac{N_{\phi}}{2}}(V^{\text{con}}_{k,p}b^{\dagger}_{q}b^{\dagger}_{q+k+p}b_{q+k}b_{q+p}+\text{h.c.}).$ (1)
@@
The effective model is one-dimensional and consists of squeezing terms in momentum space, that conserve both the total charge $\tilde{\mathcal{N}}=\sum_{q}b^{\dagger}_{q}b_{q}$ as well as the dipole moment in the LLL $\tilde{\mathcal{P}}=\sum_{q}q\,b^{\dagger}_{q}b_{q}$; Fig. 1(c). The squeezing motion can be understood from the conservation of total momenta in the $y$-direction, since the momentum labels the sites of the one-dimensional lattice. When assuming contact interactions in the unprojected system and considering the LLL, the matrix elements of the projected potential are $V^{\text{con}}_{k,p}\propto\exp[{-\frac{k^{2}+p^{2}}{2\sigma^{2}}}]/[(1+\delta_{p,0})(1+\delta_{k,{N_{y}}/{2}})]$, where the typical length scale $\sigma=L_{y}/{2}\pi l_{b}$ and the denominator avoids double counting. The effective matrix elements $V^{\text{con}}_{k,m}$ of the projected Hamiltonian can be tuned by changing the magnetic length $l_{b}$, or equivalently the thickness of the cylinder, see Fig. 1 (d). As $l_{b}$ increases the elements of the potential decay faster, and in turn a slower dynamics in the LLL arises. This is directly obtained from the dependence of the single-particle eigenstates on the magnetic length.
@@
An equivalent effective model can be obtained for interacting bosons in a 2D lattice with magnetic field, described by the Harper-Hofstadter model
@@
The non-interacting system $H_{0}$ is invariant under magnetic translations,
@@
[…]
as $[H_{0},T_{y}]=[H_{0},T_{x}]=0$; since moreover $[T_{x}^{Q},T_{y}]=0$, the three operators $H_{0},T_{x}^{Q}$and $T_{y}$ can be diagonalized simultaneously. In contrast to the continuum, the bands of the Hubbard-Hofstadter model are not flat for generic parameters. This leads to additional terms in the effective Hamiltonian projected onto the lowest band compared to Eq. (1). Interestingly, it has been found that it is possible to define geometric constraints on the lattice such that for finite-size systems the resulting bands are flat, effectively one-dimensional, and topologically non-trivial [41]. Due to the finite system this result does not contradict certain no-go theorems on the absence of flat topological bands in lattice systems with locality [42].
Non-trivial, one-dimensional flat bands are obtained when the lattice has $N_{y}$ sites in the $y$-direction and $N_{x}=Q$ sites in the $x$-direction with $Q=\alpha N_{y}$, and $\alpha$ integer. Then the form of the eigenfunctions is $\psi_{K_{y},n}(x,y)=u_{K_{y},n}(x)e^{-i\frac{2\pi}{N_{y}}K_{y}y}/\sqrt{(}N_{y})$ where $K_{y}\in[0,...N_{y}-1]$ are quantum numbers of the operator $T_{y}$, $n$ defines the energy level and $u_{K_{y},n}(x)$ are periodic functions in the $x$-direction $u_{K_{y},n}(x+Q)=u_{K_{y},n}(x)$. For these parameters the Hofstadter equations
map onto themselves, guaranteeing flat bands [41]. Similarly to the continuum case, one can interpret each flat band as a one-dimensional lattice whose sites are identified by the quantum numbers $K_{y}$. Moreover, due to the lattice geometry the functions $u_{K_{y},n}(x)$ for different $K_{y}\not=K_{y}^{\prime}$ are the same apart from a translation in the $x$-direction. Projecting the on-site interaction with a projector $P_{0}$ onto the LLL, again leads to an effective one-dimensional dipole-conserving model
[…]
We present the details of this calculation in App. A. The matrix elements $V_{k,p}^{\text{lat}}\sim\exp[{-\frac{k^{2}+p^{2}}{2\sigma^{2}}}]/[(1+\delta_{p,0})(1+\delta_{k,{N_{y}}/{2}})]$ are approximately scaling as a Gaussian, similarly to the continuum, and are controlled by the width of the single-particle orbitals that are tunable by the ratio of the hopping matrix elements $t_{y}/t_{x}$; see Fig. 1(e). The dipole squeezing motion is again suppressed with increasing overall separation of the hopping processes via a Gaussian shaped potential $V_{k,m}$. Consequences of relaxing the flat-band conditions are discussed in App. B.
@@ III Dynamics in the Lowest Landau Level
For concreteness, we focus here on the projected Hubbard-Hofstadter model. The results apply similarly for the continuum model, as only the matrix elements of the squeezing terms in Eq. (1) and Eq. (5) slightly differ. The matrix elements are determined by the typical length scale $\sigma$ of the Gaussian envelop function. For $\sigma\lesssim 1$ the so-called thin torus limit is realized and the dominant contribution to the Hamiltonian are on the one hand the density-density interactions ${n}_{q}{n}_{q+1}$, where ${n}_{q}=b^{\dagger}_{q}b_{q}$ and on the other hand the shortest-distance squeezing term $b_{q-1}^{\dagger}b^{\dagger}_{q}b_{q+1}b_{q+2}$. When considering only this contribution and a maximum local occupation of one boson per site, the system exhibits strong Hilbert space fragmentation: even when fixing the charge and dipole quantum number, the Hilbert space fragments into exponentially many sectors, each of which is parametrically small compared to the total dimension of the Hilbert space sector [43, 44, 45, 46]. As a consequence, correlation functions do not decay in time. Longer-range terms in the Hamiltonian give rise to weak Hilbert space fragmentation and restore ergodicity; for a given charge and dipole density an excitation relaxes [43]. Even though contributions to the Hamiltonian are exponentially suppressed at larger distances, they still exist and thus lead, in the thermodynamic limit, to ergodic relaxation dynamics at a certain time scale.
@@
Our goal is now to study this crossover at finite times by tuning the typical interaction length scale $\sigma$. To this end, we numerically compute the time evolution of the system in an infinite temperature state of the LLL using exact diagonalization. The initial state is thus given by the ensemble where all the states in the sector appear with the same probability. As the evolution of the full ensemble is computationally intractable for large systems, we sample product states from the respective Hilbert space sector with fixed particle density of $\nu=\tilde{\mathcal{N}}/N_{y}=1/2$ and dipole moment $\tilde{\mathcal{P}}=0$ (modulo $N_{y}$), corresponding to the largest sector. This sector contains one of the two equivalent charge density wave states. In fact, in the thin-torus limit, $\sigma\to 0$, these charge density wave states correspond to the bosonic FQH ground states at $\nu=1/2$ [30]. To set the matrix elements $V^{\text{lat}}_{k,p}$, we fix the ratio $\alpha=Q/N_{y}=2$ and vary the effective length scale $\sigma$ by tuning $t_{y}/t_{x}$. We then compute the autocorrelation function of the density $C^{0}(t)=\langle\Delta n_{0}(t)\Delta n_{0}(0)\rangle$, where $\Delta n_{0}=b^{\dagger}_{0}b_{0}-\langle b^{\dagger}_{0}b_{0}\rangle{}_{\rho_{0}}$ and $\langle...\rangle{}_{\rho_{0}}$ is the average over the sector of the density at the center of the lattice; see Fig. 2. The data is obtained for $N_{y}=22$; finite size effects are discussed in App. C. After some transient dynamics, the autocorrelation function approaches a powerlaw for late times in the thick torus limit of large $\sigma\gg 1$ with exponent consistent with $1/4$. Decreasing $\sigma$, the dynamics remains stuck on the accessible time scales indicating a prethermal Hilbert-space fragmentation regime [43].
The late-time relaxation dynamics is solely determined by the conservation laws and follows a hydrodynamics description. For a dipole-conserving systems, conventional diffusive hydrodynamics does not hold because of the restricted mobility of the system. In particular, the mobile objects are dipoles while single particles cannot move [39]. From Noether’s theorem for particle number conservation one has $\partial_{t}\rho=-\partial_{x}J$, where $\rho$ is the coarse grained density and $J$ is the particle current on the effective one-dimensional lattice. Since the mobile objects are dipoles, a finite charge current $J$ is obtained when dipoles are moving. The particle current $J$ is related to the dipole current $J_{d}$ as $J=-\partial_{x}J_{d}$, as dipoles can be interpreted as particle-hole composites [47, 48, 38]. As a next step, hydrodynamic assumptions allow us to express the dipole current in terms of density derivatives. In particular, as the dipole current is invariant under spatial inversion, the power of the spatial derivative needs to be even, leading to the lowest-symmetry allowed expansion $J_{d}=-D\partial_{x}^{2}\rho$ [39, 49]. Thus, one obtains the hydrodynamics equation for dipole-conserving systems [39, 38]
@@
In the presence of long-range, powerlaw decaying couplings, recently modifications of this hydrodynamic description have been derived [50, 51, 52]. In our case, the coefficients of the projected Hamiltonian, however, decay exponentially which leads the effective fracton hydrodynamics of Eq. (6) unchanged. We now write the evolution of the local density as
@@
We evaluate the fluctuations numerically following a quantum quench of a product initial state. The bipartite number fluctuations $\Delta^{2}N$ grow with time and saturate to a finite value due to finite size effects; Fig. 3 (a). We can best extract the hydrodynamic relaxation dynamics by directly computing the deviations of the fluctuations from their equilibrium value; Fig. 3 (b). Similarly to the autocorrelation function, the bipartite fluctuations show hydrodynamic scaling at late times for the thick torus limit $\sigma\gg 1$, while they are still stuck on the accessible scales in the thin torus limit $\sigma\lesssim 1$. We observe that the hydrodynamic behavior in the fluctuations sets in at similar times compared to the autocorrelation. This is in agreement with the hydrodynamic analysis, see solution of Eq. (6) and Eq. (9).

### v2

# Emergent Fracton Hydrodynamics of Ultracold Atoms in Partially Filled Landau Levels

@@ I Introduction
In solid state systems, fractional quantum Hall states typically arise under extreme conditions of ultra-low temperatures, large magnetic fields, and strong interactions [1, 2]. These exotic states of matter with topological order exhibit charge fractionalization and anyonic excitations. Recent experimental and technical advances enabled the study of quantum Hall physics with synthetic quantum systems as well, which offer new routes to probe these exotic states. Synthetic systems of ultracold atoms are charge neutral and hence do not minimally couple to a magnetic field, instead artificial gauge fields have to be engineered. One way to realize artificial gauge fields for ultracold quantum gases in the continuum is by using the equivalence of neutral particles under rotation and charged particles in a magnetic field [3, 4, 5, 6, 7]. By properly designing the rotating trapping potential, recent work has demonstrated the geometric squeezing of a condensate into the Lowest Landau Level (LLL) [8, 9, 10, 11]. Moreover, few-body Laughlin states have been created by rotating fermions in an optical tweezer [12]. Artificial magnetic fields have also been realized for ultracold quantum gases by spatially dependent optical couplings [13, 14, 15] and in optical lattices by laser dressing [16, 17, 18, 19, 20, 21, 22, 23, 24]. The latter approach enabled the observation of interaction-induced propagation of chiral excitations [25] as well as the preparation of fractional quantum Hall (FQH) states in the few-body limit [26]. Furthermore, FQH states have been investigated also with photons in circuit QED and cavities [27, 28, 29].
Key experimental challenges in the field are to realize the Hamiltonian with synthetic gauge fields, to cool the system to the LLL with a finite density of particles, and to prepare the FQH ground state. In this work, we propose to study the non-equilibrium relaxation dynamics to identify key signatures of the partially filled LLL at finite energy densities. Recently, it has been shown that quenching the effective mass tensor excites the dipole and graviton modes of FQH states [30, 31, 32]. These quench protocols are initialized with the ground state of the system. By contrast, here, we will avoid the challenge of preparing the interacting ground state and consider relaxation at finite energy densities.
@@
It is the common anticipation that generic quantum many-body systems at finite energy densities relax diffusively in time. How quantum systems quenched to the LL relax is an interesting question, as they are subjected to strong constraints. In this work, we show that the relaxation dynamics is strongly modified in these systems and becomes subdiffusive. In particular, we explore the unique signatures of fractionally occupied states in the LLL by investigating the relaxation of density excitations in the far-from equilibrium dynamics of a finite energy-density state in the LLL. It is known that interacting particles in the LLL are described by effective models which conserve both the global charge as well as the associated center of mass (or equivalently the dipole moment) [33, 34, 35, 36, 37, 38, 39, 40, 41, 42, 36, 43]. Due to these effective conservation laws, the mobility of the particles is constrained. Following the predictions of fracton hydrodynamics, that emerges in systems with both charge and dipole conservation [44, 45, 46], we show that these systems exhibit slow relaxation dynamics due to interaction-induced intra-Landau-level scattering processes. This unconventional relaxation dynamics thus characterizes the partially-filled LLL, which can be directly probed in experimental platforms of ultracold atoms.
@@
Our results are organized as follows. In Sec. II we introduce the effective model for the LLL; in Sec. III we address effective high-temperature properties of states in the LLL by analyzing the emergent fracton hydrodynamics of interacting particles; we find anomalous relaxation dynamics of the density-density autocorrelation function as well as the bipartite fluctuations of the particle number in a subregion of the system. In Sec. IV we discuss experimental platforms for realizing FQH physics with ultracold atoms, focusing both on rotating quantum gases and on laser-dressed atoms in optical lattices. In both cases, we analyze the full, unprojected dynamics and determine the conditions for conserving the dipole moment in real space. An outlook and discussions are provided in Sec. V.
@@ II Effective model for the Lowest Landau Level
Figure 1: Projection onto the lowest Landau level. (a) Illustration of the continuum model on a cylinder. (b) Illustration of the Harper-Hofstadter model on a torus in the Landau gauge; flat bands are realized for specific lattice geometries, as discussed in the main text. (c) When projecting onto the lowest band, the contact interaction gives rise to center-of-mass (i.e., dipole-moment) conserving hopping terms, whose matrix elements decay with typical length scale $\sigma$. (d) For the continuum model the typical length scale $\sigma$ depends only on the ratio $L_{y}/l_{b}$. (e) For the interacting Harper-Hofstadter model $\sigma$ can be tuned by changing the hopping ratio $t_{y}/t_{x}$ ($\sigma$ is numerically evaluated for system size of $N_{x}=44$ and $N_{y}=22$).
We will now consider particles confined to a two-dimensional plane subjected to a strong perpendicular magnetic field, such that Landau levels form. When considering a finite energy-density state within the LLL, their dynamics remains restricted to the LLL when the interaction energy between particles is small compared to the energy gap between Landau levels. Consequently, relaxation is governed by interactions projected onto the LLL. As the matrix elements of the projected interactions introduce a notion of distance between different quantum numbers, the relaxation effectively occurs in a one dimensional lattice. We will now derive the projected models both in the continuum and on a lattice. In both cases, we demonstrate that the system possesses an emergent dipole-conserving dynamics within this effective one-dimensional lattice.
@@
We first focus on the continuum, where the single-particle Hamiltonian that describes the system is $H_{0}=\frac{1}{2m}(\mathbb{p}-\mathbb{A})^{2}$. Here, $(\mathbb{p}-\mathbb{A})$ is the generalized momentum in the presence of the magnetic field, and $\mathbb{A}=B\,(0,x)$ is the vector potential in the Landau gauge. The properties of the system are determined by the magnetic length $l_{b}=\sqrt{\frac{\hbar}{eB}}$, the number of magnetic fluxes $N_{\phi}=\frac{L_{x}L_{y}}{2\pi l^{2}_{b}}$, and the cyclotron frequency $\omega_{c}=\frac{eB}{m}$. Assuming an infinite cylinder geometry with periodic boundary conditions in the $y$-direction, Fig. 1(a), the basis that diagonalizes the single-particle Hamiltonian is [33, 34, 35, 42], $\psi_{\gamma}^{m}(\mathbb{r})=\frac{1}{\sqrt{{\pi}\gamma!2^{\gamma}L_{y}}}H_{\gamma}\bigg{(}\frac{x-x_{m}}{l_{B}}\bigg{)}e^{-\frac{1}{2}(\frac{x-x_{m}}{l_{b}})^{2}}e^{ik_{m}y}$, where $H_{n}$ are the Hermite polynomials, $x_{m}=2\pi m\frac{l^{2}_{b}}{L_{y}}$ with $m\in[0,N_{\phi}-1]$ determine the center of the wave function, and $k_{m}=\frac{2\pi m}{L_{y}}$ is the momentum in $y$-direction. The associated single-particle eigenvalues are $E_{\gamma}^{m}=\hbar\omega_{c}(\gamma+\frac{1}{2})$. Assuming that particles initially occupy only one Landau level, density-density interactions $V=\sum_{i,j}V_{i,j}n_{i}n_{j}$ of magnitude much smaller than the inter-level spacing $\omega_{c}$, give rise to an effective dynamics that is confined to the same Landau level. The Hamiltonian is obtained by projecting the interacting Hamiltonian $H=H_{0}+V$ onto the Landau level [42]
$P_{0}HP_{0}=\sum_{q=0}^{N_{\phi}-1}\sum_{p=0}^{\frac{N_{\phi}}{2}-1}\sum_{k=p+1}^{\frac{N_{\phi}}{2}}(V^{\text{con}}_{k,p}b^{\dagger}_{q}b^{\dagger}_{q+k+p}b_{q+k}b_{q+p}+\text{h.c.}),$ (1)
@@
where $p,q,k\,\in[0,N_{\phi}-1]$ are integers that label the $y-$component of the momenta in the LLL, and $N_{\Phi}$ coincide with the number of magnetic cells in the $y-$direction. The effective model is one-dimensional and consists of squeezing terms in momentum space, that conserve both the total charge $\tilde{\mathcal{N}}=\sum_{q}b^{\dagger}_{q}b_{q}$ as well as the dipole moment in the LLL $\tilde{\mathcal{P}}=\sum_{q}q\,b^{\dagger}_{q}b_{q}$; Fig. 1(c). The squeezing motion can be understood from the conservation of total momenta in the $y$-direction, since the momentum labels the sites of the one-dimensional lattice. When assuming contact interactions in the unprojected system and considering the LLL, the matrix elements of the projected interaction depends only on the overlap in the $x$-component of the eigenstates in the LLL. As such, they can be tuned by changing the parameters in the Hamiltonian. For our model, $V^{\text{con}}_{k,p}\propto\exp[{-\frac{k^{2}+p^{2}}{2\sigma^{2}}}]/[(1+\delta_{p,0})(1+\delta_{k,{N_{y}}/{2}})]$ [42, 47], where the typical length scale $\sigma=L_{y}/{2}\pi l_{b}$ and the denominator avoids double counting. The effective matrix elements $V^{\text{con}}_{k,m}$ of the projected Hamiltonian can be tuned by changing the magnetic length $l_{b}$, or equivalently the thickness of the cylinder, see Fig. 1 (d). As $l_{b}$ increases the elements of the potential decay faster, and in turn a slower dynamics in the LLL arises. This is directly obtained from the dependence of the single-particle eigenstates on the magnetic length.
@@
An equivalent effective model can be obtained for interacting bosons in a 2D lattice with magnetic field, described by the interacting Harper-Hofstadter model
@@
where $n_{\phi}=P/Q$ is the number of fluxes per unit cell, $P,Q$ integers and coprime. The non-interacting system $H_{0}$ is invariant under magnetic translations,
$T_{x}=\sum_{x,y}e^{i2\pi n_{\phi}y}a^{\dagger}_{x+1,y}a_{x,y}$
$T_{y}=\sum_{x,y}a^{\dagger}_{x,y+1}a_{x,y},$ (3)
as $[H_{0},T_{y}]=[H_{0},T_{x}]=0$; since moreover $[T_{x}^{Q},T_{y}]=0$, the three operators $H_{0},T_{x}^{Q}$and $T_{y}$ can be diagonalized simultaneously. In contrast to the continuum, the bands of the Harper-Hofstadter model for $U=0$ are not flat for generic parameters. This leads to additional terms in the effective Hamiltonian projected onto the lowest band compared to Eq. (1). Interestingly, it has been found that it is possible to define geometric constraints on the lattice such that for finite-size systems the resulting bands are flat, effectively one-dimensional, and topologically non-trivial [48]. Due to the finite system this result does not contradict certain no-go theorems on the absence of flat topological bands in lattice systems with locality [49].
Non-trivial, one-dimensional flat bands are obtained when the lattice has $N_{x}=Q$ sites in the $x$-direction with $Q=\alpha N_{y}$, $\alpha$ integer, and $N_{y}$ sites in the $y$-direction, that in this direction corresponds also to the number of magnetic unit cells. Then the form of the eigenfunctions is $\psi_{K_{y},n}(x,y)=u_{K_{y},n}(x)e^{-i\frac{2\pi}{N_{y}}K_{y}y}/\sqrt{(}N_{y})$ where $K_{y}\in[0,...N_{y}-1]$ are quantum numbers of the operator $T_{y}$, $n$ defines the energy level and $u_{K_{y},n}(x)$ are periodic functions in the $x$-direction $u_{K_{y},n}(x+Q)=u_{K_{y},n}(x)$. For these parameters the Hofstadter equations
are the same as the momenta are changed, up to a translation in the $x-$direction, guaranteeing flat bands [48]. Similarly to the continuum case, one can interpret each flat band as a one-dimensional lattice whose sites are identified by the quantum numbers $K_{y}$. Moreover, due to the lattice geometry the functions $u_{K_{y},n}(x)$ for different $K_{y}\not=K_{y}^{\prime}$ are the same apart from a translation in the $x$-direction. Projecting the on-site interaction with a projector $P_{0}$ onto the LLL, again leads to an effective one-dimensional dipole-conserving model
@@
We present the details of this calculation in App. A. The matrix elements $V_{k,p}^{\text{lat}}\sim\exp[{-\frac{k^{2}+p^{2}}{2\sigma^{2}}}]/[(1+\delta_{p,0})(1+\delta_{k,{N_{y}}/{2}})]$ are approximately scaling as a Gaussian, similarly to the continuum, and are controlled by the width of the single-particle orbitals that are tunable by the ratio of the hopping matrix elements $t_{y}/t_{x}$; see Fig. 1(e). The dipole squeezing motion is again suppressed with increasing overall separation of the hopping processes via a Gaussian shaped potential $V_{k,m}$. Even thought the precise geometric constraints may be challenging to realize experimentally, this flat-band model severs as a controlled starting point for our analysis. Consequences of relaxing the flat-band conditions are discussed in App. B. From this model, the continuum limit is recovered by sending $Q\to\infty$; in this limit, up to exponentially small corrections, the number of sites in the effective one-dimensional model Eq. (5) corresponds to the number of magnetic fluxes [48, 50, 51, 52].
Thus, for both models analyzed in this section, we have obtained effective dipole conserving dynamics in a one dimensional lattice. For both cases, the mechanism that allows for the emerging dipole conservation in the LLL is the conservation of one of the components of momentum. We will now show how this unconventional conservation law determines subdiffusive behavior in the LLL and how the geometry of the model can be used to tune dynamical properties.
@@ III Dynamics in the Lowest Landau Level
For concreteness, we focus here on the projected Harper-Hofstadter model. The results apply similarly for the continuum model, as only the matrix elements of the squeezing terms in Eq. (1) and Eq. (5) slightly differ. The matrix elements are determined by the typical length scale $\sigma$ of the Gaussian envelop function. For $\sigma\lesssim 1$ the so-called thin torus limit is realized and the dominant contribution to the Hamiltonian are on the one hand the density-density interactions ${n}_{q}{n}_{q+1}$, where ${n}_{q}=b^{\dagger}_{q}b_{q}$ and on the other hand the shortest-distance squeezing term $b_{q-1}^{\dagger}b_{q}b_{q+1}b^{\dagger}_{q+2}$. When considering only this contribution and a maximum local occupation of one boson per site, the system exhibits strong Hilbert space fragmentation: even when fixing the charge and dipole quantum number, the Hilbert space fragments into exponentially many Krylov subspaces, each of which is parametrically small compared to the total dimension of the Hilbert space sector [53, 54, 55, 56]. As a consequence, correlation functions do not decay in time. Longer-range terms in the Hamiltonian give rise to weak Hilbert space fragmentation and restore ergodicity; for a given charge and dipole density an excitation relaxes [53]. Even though contributions to the Hamiltonian are exponentially suppressed at larger distances, they still exist and thus lead, in the thermodynamic limit, to ergodic relaxation dynamics at a certain time scale.
@@
Our goal is now to study this crossover at finite times by tuning the typical interaction length scale $\sigma$. We numerically compute the time evolution of the system in an infinite temperature state of the LLL using exact diagonalization. The initial state serves as a generic finite energy-density state in the LLL. While the short-time dynamics depends on the details of the initial state, the emergence of hydrodynamics relies only on the symmetries of the problem. Similar late-time dynamics to the one presented here, will therefore also arise when a single mode of the LLL is highly occupied [8] or for other experimentally relevant initial states that are restricted to the LLL. In order to numerically compute the correlation functions of the high-temperature state, we sample product states from the respective Hilbert space sector with fixed particle density of $\nu=\tilde{\mathcal{N}}/N_{y}=1/2$ and dipole moment $\tilde{\mathcal{P}}=0$ (modulo $N_{y}$), corresponding to the largest sector. This sector contains one of the two equivalent charge density wave states. In fact, in the thin-torus limit, $\sigma\sim 0$, these charge density wave states correspond to the bosonic FQH ground states at $\nu=1/2$ [36]. To set the matrix elements $V^{\text{lat}}_{k,p}$, we fix the ratio $\alpha=Q/N_{y}=2$ and vary the effective length scale $\sigma$ by tuning $t_{y}/t_{x}$. We then compute the autocorrelation function of the density $C^{0}(t)=\langle\Delta n_{0}(t)\Delta n_{0}(0)\rangle$, where $\Delta n_{0}=b^{\dagger}_{0}b_{0}-\langle b^{\dagger}_{0}b_{0}\rangle{}_{\rho_{0}}$ and $\langle...\rangle{}_{\rho_{0}}$ is the average over the sector; see Fig. 2. The data is obtained for $N_{y}=22$; finite size effects are discussed in App. C. After some state-dependent transient dynamics, the autocorrelation function approaches a powerlaw for late times in the thick torus limit of large $\sigma\gtrsim 1$ with exponent of $\sim 1/4$. As will be shown later, this result is consistent with the late-time behavior of a dipole-moment conserving system. Decreasing $\sigma$, the dynamics remains stuck on the accessible time scales indicating a prethermal-regime of strong Hilbert-space fragmentation [53].
[…]
As a next step, hydrodynamic assumptions allow us to express the dipole current in terms of density derivatives. In particular, as the dipole current is invariant under spatial inversion, the power of the spatial derivative needs to be even, leading to the lowest-symmetry allowed expansion $J_{d}=-D\partial_{x}^{2}\rho$ [45, 59]. Thus, one obtains the hydrodynamics equation for dipole-conserving systems [45, 44]
@@
In the presence of long-range, powerlaw decaying couplings, recently modifications of this hydrodynamic description have been derived [60, 61, 62]. In our case, the coefficients of the projected Hamiltonian, however, decay exponentially which leads the effective fracton hydrodynamics of Eq. (6) unchanged.
We now write the evolution of the local density as
@@
Thus the fluctuations approach their finite-size saturation value hydrodynamically.
We evaluate the fluctuations numerically following a quantum quench of a product initial state. The bipartite number fluctuations $\Delta^{2}N$ grow with time and saturate to a finite value due to the finite system size; Fig. 3 (a). We can best extract the hydrodynamic relaxation dynamics by directly computing the deviations of the fluctuations from their late-time saturation value; Fig. 3 (b). We observe that similarly to the autocorrelation function, the hydrodynamic behavior in the fluctuations sets in at later times as the interaction length scale $\sigma$ is decreased. For $\sigma=1.4$ the onset time of hydrodynamics is difficult to estimate due to the very slow relaxation in the non-universal regime. For sufficiently small $\sigma\lesssim 1$, the hydrodynamic regime is not observable on the time scales and system sizes accessible in our numerical simulations; as the dominant terms in the model are of shortest possible range giving rise to strong Hilbert space fragmentation with exponentially many disconnected Krylov subspaces [53]. Nonetheless, due to the Gaussian decaying coupling terms, strong Hilbert space fragmentation is always formally avoided in our model for any finite interaction length scale $\sigma$, implying that in the thermodynamic limit and at late times the system will ultimately be able to thermalize.